• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric Voutier

    Log entry time 00:08:52 on April 2,2000

    Entry number 37650

    keyword=Shift Summary


    New crew on shift: Gary Rutledge, Eric Voutier
    ***************

    16:00 Beam is being tuned for 5 pass setting

    19:00 Beam is close to the hall

    19:50 Make an entry into the Hall to reset Q3 in the electron arm

    20:50 While exiting the Hall, MCC experiences RF separator problems

    21:30 Q3 cycling ends, start controling OTR and beam spot size; the analysis of the
    HARP #56 scan on jlabs2 is not working; from the scanner's picture, we anticipate
    a wide spot size about 300um in both directions... A fine tuning is not needed for
    Nilanga purposes (according to his specs).

    22:00 Fast raster problems

    22:25 Moving to Carbon target

    22:30 Acquisition system not responding, ROC14 problem.
    ....... ER1 problem is also there when running on the carbon target

    23:15 Move to LH2 target

    #-- Run history
    --------

    ......... Kinematics 301

    ..3130.... 12 Carbon ......
    ..3131.... 12 Carbon ...... 7/10/1/10/1 ...... I = 10 uA
    ..3132.... LH2 ............... I = 2 uA / good run / large raster
    ..3133.... LH2 ............... I = 2 uA / good run / reduced size raster
    ..3144.... LH2 ............... I = 2 uA goes to next shift